OpenStudy (anonymous):

a)If an inter n ends in 3, show why the last digit of n^5 is 3. b)Prove that, for all integers n, the last digit of n^5 is the same as the last digit of n. c) Show that n^5n is divisible by 30 for all integers n. PLEASE HELP ITS VERY URGENT- MUST BE DONE BEFORE TUESDAY :(

OpenStudy (anonymous):

btw please giv a detailed response coz im not very smart :P

OpenStudy (anonymous):

a) You can form a kind of cycle to find the units digit of numbers like n^5. Let n be 13. 13^1 --> units digit 3 13^2 --> units digit 3*3 = 9 13^3 --> units digit = units digit of 3*3*3 = units digit of 27 = 7

OpenStudy (anonymous):

Keep repeating this till 13^5

OpenStudy (anonymous):

Units digit of 13^5 = units digit of 3^5 = units digit of 3*3*3*3*3

OpenStudy (anonymous):

Do you know how to do it now?

OpenStudy (anonymous):

Are you there @cloe17

OpenStudy (anonymous):

oops soz im here now

OpenStudy (anonymous):

Did you understand my explanation?

OpenStudy (anonymous):

but does that show that every number ending in 3^5 will end in 3?

OpenStudy (anonymous):

You're asking for the next question right? You can try with any number for second question but that doesn't mean proving. I still don't know how to PROVE the second one.

OpenStudy (anonymous):

Did you get the first one?

OpenStudy (anonymous):

alternate way, if you have done binomial expansion: if \[n = 10k + 3\] \[n^5 = (10k + 3)^5 =243+4050 k+27000 k^2+90000 k^3+150000 k^4+100000 k^5\] \[n^5 = 10(24+405 k+2700 k^2+9000 k^3+15000 k^4+10000 k^5) + 3\]

OpenStudy (anonymous):

for the second one you could let n = 10k + r

OpenStudy (anonymous):

That's the format for a two digit number.

OpenStudy (anonymous):

no its not, i could let k = 15

OpenStudy (anonymous):

and get 150 + r

OpenStudy (anonymous):

I understand your expansion but how does that help us in proving units digit is n?

OpenStudy (anonymous):

well that shows the last digit is three if thats what you mean, because when we divide by 10 we are left with a remainder of 3

OpenStudy (anonymous):

Oh yes!! Hope that clears your doubts @cloe17

OpenStudy (anonymous):

...

OpenStudy (anonymous):

i dont get why n=10k + 3?

OpenStudy (anonymous):

this way sorts out a) and b) together: consider two positive integers, n and m the last digit of n is p , and the last digit of m is q so we have n=10a + p m = 10b + q where a and b are integers. now lets try and find the last digit of nm nm = (10a+p)(10b+q) = 100ab + 10aq+ 10pb + pq = 10(10ab + aq+ pb ) + pq so therefore the last digit of mn is the last digit of pq eg lets find the last digit of 1412359 x 458789343 well we just need to find the last digit of 9 x 3 9 x 3 = 21 so the last digit of 1412359 x 458789343 is the last digit of 21, which is 1

OpenStudy (anonymous):

would you agree we can write any number ending in 3 as 10k + 3 ? eg 23 = 10(2) + 3 153 = 10(15) +3 34873843 = 10(3487384) + 3 so we can say that if n ends in 3, then we can write it as 10k +3

OpenStudy (anonymous):

sorry i still dont really gwt it, im only in yr 7 so dont worry :D

OpenStudy (anonymous):

can you just tell me how to do c?

OpenStudy (anonymous):

wait actually i get the 10k + 3 formula now except im not sure why k has to be multiplied bt 10

OpenStudy (anonymous):

oh ok sorry, i thought you were older these are pretty tricky for year 7 im impressed c) Show that n^5n is divisible by 30 for all integers n do you mean n^(5n) ? because that doesnt work for n = 1 for example

OpenStudy (anonymous):

n^5 - n

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

try to make the answer my level is that okay?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

i'm struggling to explain it simpler.. can i try and guide you through the answer i think is the simplest and easiest?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

looking at each part of the problem gives us hints for the next part ok so for a) we need to think about numbers that end in 3 every number ending in 3 can be split up like this:  would you agree?